# Borsuk-Ulam and Replicable Learning of Large-Margin Halfspaces

### v3

@@ 1.2 Concluding remarks and open problems
While Theorem 1.4 provides an elegant and comprehensive characterization of various notions of replicability and privacy for total concepts through the combinatorial framework of the Littlestone dimension, the landscape of replicability for partial classes, as demonstrated by the results of this paper, is more intricate and less understood.
The “DP-learnability to Shared-randomness Replicability” reduction from [BGH+23] extends to the partial setting. Moreover, [FHM+24] recently showed that for partial classes, DP-learnability implies a finite Littlestone dimension.
Theorem 1.16 ([BGH+23, FHM+24]).
Let {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ be a partial concept class.
  • If $\mathcal{H}$ is DP-learnable, then $\mathcal{H}$ is shared-randomness replicable.
  • If $\mathcal{H}$ is DP-learnable, then $\operatorname{Ldim}(\mathcal{H})<\infty$.
On the other hand, our main theorem shows that for partial concepts, list replicability does not follow from bounded Littlestone dimension, shared-randomness replicability, DP-learnability, or even pure DP-learnability. It is also straightforward to show that, even in the partial setting, list replicability implies shared-randomness replicability.
Proposition 1.17 ([KKMV23, Lemma 8]).
If {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is list replicable (equiv. globally stable), then $\mathcal{H}$ is shared-randomness replicable.
To our knowledge, no further relationships among the four notions of DP-learnability, shared-randomness replicability, Littlestone dimension, and list replicability are known for partial functions.
Open problems.
@@
    Is there a combinatorial notion of dimension that characterizes list replicability?
@@
    $\operatorname{VCdim}(\mathcal{H})\leq\widetilde{\operatorname{LR}}(\mathcal{H})\leq\operatorname{rank_{\pm}}(\mathcal{H})?$
@@
    In an article posted on arXiv just a few days ago, Chornomaz, Moran, and Waknine [CMW25] explored this problem using a topological approach, but the question remains open.
2 Preliminaries
Much of the notation and definitions were already outlined in the introduction. In this section, we review a few key concepts from learning theory, specifically PAC learnability, the VC and Littlestone dimensions, and differential privacy.
In some of our proofs, we will apply Hoeffding’s classic concentration inequality, which we state below for reference.
Theorem 2.1 (Hoeffding’s inequality).
Let $c\in\mathbb{R}$ and let $\bm{x}_{1},\ldots,\bm{x}_{n}$ be independent random variables with $\bm{x}_{i}\in[-c,c]$ and $\mathbb{E}[\bm{x}_{i}]=0$. Then, for any $t>0$,
$\Pr\left[\left|\sum_{i=1}^{n}\bm{x}_{i}\right|\geq t\right]\leq 2e^{-\frac{t^{2}}{2c}}.$
We define the population loss of a partial concept $h\in\left\{\pm 1,\star\right\}^{\mathcal{X}}$ with respect to a distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm 1\right\}$ as
$\mathcal{L}_{\mathcal{D}}(h)\coloneqq\Pr_{(\bm{x},\bm{y})\sim\mathcal{D}}[h(\bm{x})\neq\bm{y}].$
Under this definition, predictions with $h(x)=\star$ always count as mistakes. The population loss of a class $\mathcal{H}$ is given by the smallest population loss achievable within $\mathcal{H}$:
$\mathcal{L}_{\mathcal{D}}(\mathcal{H})\coloneqq\inf_{h\in\mathcal{H}}\mathcal{L}_{\mathcal{D}}(h).$
The following simple lemma from [AHHM22] establishes the connection between realizability and having zero population loss.
Lemma 2.2 ([AHHM22]).
Let {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ be a partial concept class, and let $\mathcal{D}$ be a distribution on $\mathcal{X}\times\left\{\pm 1\right\}$. If $\mathcal{L}_{\mathcal{D}}(\mathcal{H})=0$, then $\mathcal{D}$ is realizable by $\mathcal{H}$. Conversely, if $\mathcal{D}$ is realizable and has finite or countable support, then $\mathcal{L}_{\mathcal{D}}(\mathcal{H})=0$.
PAC learning of partial classes:
Given a concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$, the goal of PAC (Probably Approximately Correct) learning is for the learning rule to produce, with high probability, a hypothesis whose population loss is close to the best achievable loss within $\mathcal{H}$.
Definition 2.3 (PAC learning of partial concept classes).
A class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is PAC learnable if there is a learning rule $\bm{\mathcal{A}}$ and a function $n(\epsilon,\delta)$ such that for every $\epsilon,\delta>0$ and every realizable distribution $\mathcal{D}$ on $\mathcal{X}\times\left\{\pm\right\}$, we have
$\Pr_{\bm{S}\sim\mathcal{D}^{n}}[\mathcal{L}_{\mathcal{D}}(\bm{\mathcal{A}}(\bm{S}))\leq\epsilon]\geq 1-\delta\ \text{ where }n=n(\epsilon,\delta).$
2.1 VC and Littlestone dimensions
We now extend the definition of the VC dimension to partial concept classes.
Definition 2.4 (Shattered set).
A finite set of points $C=\left\{x_{1},\ldots,x_{n}\right\}\subseteq\mathcal{X}$ is shattered by a partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ if, for every pattern $y\in\left\{\pm 1\right\}^{n}$, there exists $h\in\mathcal{H}$ such that $h(x_{i})=y_{i}$ for all $i\in[n]$.
Definition 2.5 (VC dimension).
The VC dimension of a partial concept class $\mathcal{H}$, denoted $\operatorname{VCdim}(\mathcal{H})$, is the largest integer $d$ for which there exists a size-$d$ subset of $\mathcal{X}$ that is shattered by $\mathcal{H}$. If no such maximum $d$ exists, we define $\operatorname{VCdim}(\mathcal{H})=\infty$.
Viewed as a matrix, the VC dimension of $\mathcal{H}$ is the largest $d$ such that the associated partial matrix $M_{\mathcal{X}\times\mathcal{H}}$ contains a $d\times 2^{d}$ submatrix with ±$\pm 1$ entries, where the columns enumerate all $d$-bit ±$\pm 1$ patterns.
The fundamental theorem of PAC learning asserts that a total concept class is PAC learnable if and only if its VC dimension is finite. This result was extended to partial concept classes by [AHHM22] using a completely different proof.
Theorem 2.6 ([AHHM22]).
The following statements are equivalent for any partial concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$.
  • $\operatorname{VCdim}(\mathcal{H})<\infty$;
  • $\mathcal{H}$ is PAC learnable;
Littlestone dimension:
In his influential work, Littlestone [Lit88] provided a combinatorial characterization of concept classes that are online learnable. To this end, he introduced a measure that extends the VC dimension by considering the shattering of decision trees rather than sets.
A mistake tree of depth $d$ over a domain $\mathcal{X}$ is a complete binary tree of depth $d$ with the following properties:
  • Each internal node is labeled by an element $x\in\mathcal{X}$.
  • Each edge is labeled by a binary value $b\in\left\{\pm 1\right\}$, where $b=-1$ indicates a left child and $b=1$ indicates a right child.
Every root-to-leaf path in the tree is described by a sequence
$(x_{1},b_{1}),\ldots,(x_{d},b_{d}),$
where $x_{i}\in\mathcal{X}$ is the label of the $i$th internal node along the path, and $b_{i}$ specifies whether the path moves to the left or right child at each level.
A concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ shatters a mistake tree if, for every root-to-leaf path
$(x_{1},b_{1}),\ldots,(x_{d},b_{d}),$
there exists a hypothesis $h\in\mathcal{H}$ such that $h(x_{i})=b_{i}$ for all $i\in[d]$.
Definition 2.7 (Littlestone Dimension).
The Littlestone dimension of a concept class $\mathcal{H}$, denoted $\operatorname{Ldim}(\mathcal{H})$, is the largest integer $d$ for which there exists a mistake tree of depth $d$ that is shattered by $\mathcal{H}$. If no such maximum $d$ exists, we define $\operatorname{Ldim}(\mathcal{H})=\infty$.
It always holds that $\operatorname{VCdim}(\mathcal{H})\leq\operatorname{Ldim}(\mathcal{H})$, since any set $C=\left\{x_{1},\ldots,x_{d}\right\}$ shattered by $\mathcal{H}$ gives rise to a mistake tree of depth $d$, where all nodes at level $i$ are labeled with $x_{i}$. This tree is shattered by $\mathcal{H}$.
Littlestone proved that a total concept class $\mathcal{H}$ is online learnable if and only if $\operatorname{Ldim}(\mathcal{H})<\infty$. This result was later extended to partial concept classes by [AHHM22].
2.2 Differentially Private Learnability
$p(E)\leq e^{\epsilon}q(E)+\delta\ \text{ and }\ q(E)\leq e^{\epsilon}p(E)+\delta.$
Definition 2.8 (Differential Privacy).
Given $\epsilon,\delta>0$, a randomized learning rule
In approximate differential privacy, the parameters $\epsilon$ and $\delta$ are typically set as follows: $\epsilon$ is taken to be a small fixed constant (e.g., 0.1), while $\delta$ is a negligible function, $\delta=n^{\omega(1)}$, where $n$ is the sample size. Definitions 2.9 and 2.10 formally define differentially private learnability under the assumption that the privacy parameter $\epsilon$ is fixed at $\epsilon=0.1$.
Definition 2.9 (Approximate Differentially Private Learnability).
We say that a concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is approximate differentially private learnable (DP-learnable) if there is a learning rule $\bm{\mathcal{A}}:(\mathcal{X}\times\left\{\pm 1\right\})^{*}\to\left\{\pm 1\right\}^{\mathcal{X}}$ with sample complexity $n(\epsilon,\delta,\delta^{\prime})$ such that the following holds.
  • For every $\epsilon,\delta,\delta^{\prime}>0$, the class $\mathcal{H}$ is $(\epsilon,\delta)$-PAC learnable by $\bm{\mathcal{A}}$ using $n(\epsilon,\delta,\delta^{\prime})$ samples.
  • For every $\epsilon,\delta,\delta^{\prime}>0$, the learning rule $\bm{\mathcal{A}}$ applied to samples of size $n(\epsilon,\delta,\delta^{\prime})$ is $(0.1,\delta^{\prime})$-differentially private.
  • For every fixed $\epsilon,\delta>0$, we have
    $\lim_{\delta^{\prime}\to 0}\frac{\log(1/\delta^{\prime})}{\log n(\epsilon,\delta,\delta^{\prime})}=\infty.$
Definition 2.10 (Pure Differentially Private Learnability).
We say that a concept class {±$\mathcal{H}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ is pure differentially private learnable (pure DP-learnable) if $\mathcal{H}$ is PAC learnable by a $(0.1,0)$-differentially private learning rule.
3 Proof of the Main Theorem
@@ 3.1 The Lower Bound
By the definition of list replicability, for any $\delta>0$, there is an integer $n$ so that for any realizable distribution $\mathcal{D}$, there exists a list of hypotheses $\{h_{1},\dots,h_{L}\}$ with
$\operatorname{Pr}_{\mathbf{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\mathbf{S})\in\{h_{1},\dots,h_{L}\}]\geq 1-\delta\text{ and }\mathcal{L}_{\mathcal{D}}(h_{i})\leq\epsilon\text{ for all }i\in[L].$
Now pick any $\alpha>0$ and $\epsilon^{\prime}\in(\epsilon,1/2)$. By taking $\delta$ sufficiently small, for any distribution $\mathcal{D}$, we can choose a hypothesis $h_{\mathcal{D}}\in\{h_{1},\dots,h_{L}\}$ such that
$\Pr_{\mathbf{S}\sim\mathcal{D}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{\mathcal{D}}]>\frac{1}{L+\alpha}\text{ and }\mathcal{L}_{\mathcal{D}}(h_{\mathcal{D}})<\epsilon^{\prime}.$ (3)
We will focus on a certain collection of realizable distributions $\mathcal{D}$ on $\mathbb{S}^{d-1}\times\{\pm 1\}$. For any $w\in\mathbb{S}^{d-1}$, take $\mathcal{D}_{w}$ to be the uniform distribution on the set $\{(x,h_{w}(x))\ |\ x\in{\mathrm{supp}}(h_{w})\}$. These distributions are, by definition, realizable. Hence, for each $\mathcal{D}_{w}$, we can choose some particular hypothesis $h_{\mathcal{D}_{w}}$ that satisfies the conditions in Eq. 3. Collect these hypotheses in a set $T$, that is to say
$T\coloneqq\{h_{\mathcal{D}_{w}}\ |\ w\in\mathbb{S}^{d-1}\}.$
For each $h\in T$, define the set $C_{h}\subset\mathbb{S}^{d-1}$ as
$C_{h}\coloneqq\left\{w\in\mathbb{S}^{d-1}\ |\ \Pr_{\mathbf{S}\sim\mathcal{D}_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h]>\frac{1}{L+\alpha}\text{ and }\mathcal{L}_{\mathcal{D}_{w}}(h)<\epsilon^{\prime}\right\}.$
@@
The family $\{C_{h}\}_{h\in T}$ forms an antipodal-free open cover of $\mathbb{S}^{d-1}$.
@@
The fact that any set $C_{h}$ is antipodal-free follows from the accuracy constraint $\mathcal{L}_{\mathcal{D}_{w}}(h)<\epsilon^{\prime}$. Indeed, for any $w\in\mathbb{S}^{d-1}$, the concepts $h_{w}$ and $h_{-w}$ have identical support, on which they disagree at every point. Thus the population loss of any hypothesis $h$ satisfies the equation
$\mathcal{L}_{\mathcal{D}_{w}}(h)+\mathcal{L}_{\mathcal{D}_{-w}}(h)=1.$
For any $w\in C_{h}$, we have that $\mathcal{L}_{\mathcal{D}_{w}}(h)<\epsilon^{\prime}<1/2$, whereby $w$ and $-w$ cannot both be in $C_{h}$.
Next, each set $C_{h}$ is open because both $\Pr_{\bm{S}\sim\mathcal{D}_{w}^{n}}[\bm{\mathcal{A}}(\bm{S})=h]$ and $\mathcal{L}_{\mathcal{D}_{w}}(h)$ are continuous in $w$. Lastly, the family $\{C_{h}\}_{h\in T}$ covers $\mathbb{S}^{d-1}$ because, for any $w\in\mathbb{S}^{d-1}$, the set $C_{h_{\mathcal{D}_{w}}}$ contains $w$ by construction. ∎
Now note that the antipodal-free open cover $\{C_{h}\}_{h\in T}$ admits a finite subcover by the compactness of the unit sphere. Applying Theorem 3.1 to such a finite subcover guarantees that some $w\in\mathbb{S}^{d-1}$ is contained in at least $t\coloneqq\lceil\frac{d}{2}+1\rceil$ sets $C_{h_{1}},C_{h_{2}},\dots,C_{h_{t}}$. Unpacking definitions reveals that the distribution $\mathcal{D}_{w}$ has the property
$\Pr_{\mathbf{S}\sim\mathcal{D}_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]>\frac{1}{L+\alpha}$
for $t$ distinct hypotheses $h_{i}\in T$. Because these $h_{i}$ are distinct, the events $[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]$ are disjoint, and therefore
$1\geq\operatorname{Pr}_{\mathbf{S}\sim\mathcal{D}_{w}^{n}}\bigcup_{i=1}^{t}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]=\sum_{i=1}^{t}\operatorname{Pr}_{\mathbf{S}\sim\mathcal{D}_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]>\frac{t}{L+\alpha}.$
@@ 3.2 The Upper Bound
To prove the upper bound, we design a list replicable learning algorithm $\bm{\mathcal{A}}$ that learns $\mathcal{H}_{\gamma}^{d}$ with list size $d$ independent of $\epsilon>0$. Given $w\in\mathbb{S}^{d-1}$, let $\overline{h}_{w}:\mathbb{S}^{d-1}\to\{\pm 1\}$ denote the total concept class corresponding to the closed half-space defined by $w$.
$\overline{h}_{w}(x)\coloneqq\begin{cases}1&\text{if }\langle w,x\rangle\geq 0\\ -1&\text{if }\langle w,x\rangle<0\end{cases}.$
@@
$\begin{array}[]{l rll}\max&\gamma&&\\ \text{s.t. }&y_{i}{\left\langle x_{i},w\right\rangle}&\geq\gamma&\text{for }i=1,\ldots,n\\ &\gamma&\geq 0&\\ &w&\in\mathbb{S}^{d-1}&\\ \end{array}$
@@
Furthermore, for any distribution $\mathcal{D}$ over $\mathbb{S}^{d-1}\times\{\pm 1\}$, we call $w$ a $(\gamma,\epsilon)$-separator for $\mathcal{D}$ if
$\Pr_{(x,y)\sim\mathcal{D}}[y\langle x,w\rangle<\gamma]\leq\epsilon.$
When learning $\mathcal{H}_{\gamma}^{d}$ in the realizable setting, for any sample set $S$ drawn from a realizable distribution $\mathcal{D}$, there is some $w$ that $\gamma$-separates $S$. Therefore, the hard-SVM will be feasible and return a vector $w$ that $\gamma$-separates $S$.
The following theorem, due to [STBWA98], says that if we take a sufficiently large sample $\bm{S}$ and compute a good separator $w$ for it using hard-SVM, then with high probability, $w$ will also be a good separator for $\mathcal{D}$.
@@
For all $\epsilon,\delta>0$, there exists $n\coloneqq n(\epsilon,\delta)$ such that the following holds. Let $\mathcal{D}$ be any distribution over $\mathbb{S}^{d-1}\times\left\{\pm 1\right\}$.
$\Pr_{\bm{S}\sim\mathcal{D}^{n}}\left[\text{Every $w\in\mathbb{S}^{d-1}$ that $\gamma$-separates $\bm{S}$ also $\left(\frac{\gamma}{2},\epsilon\right)$-separates $\mathcal{D}$}\right]\geq 1-\delta.$
@@
To prove Theorem 3.4, one can apply [STBWA98, Theorem 3.5] to show that, with probability at least $1-\delta$, both $h_{1}(x)\coloneqq\operatorname{sgn}({\left\langle x,w\right\rangle}+\frac{\gamma}{2})$ and $h_{2}(x)\coloneqq\operatorname{sgn}({\left\langle x,w\right\rangle}-\frac{\gamma}{2})$ have loss at most $\frac{\epsilon}{2}$, in which case $w$ is a $\left(\frac{\gamma}{2},\epsilon\right)$-separator for $\mathcal{D}$.
@@
We will also use the following simple concentration result for sums of i.i.d. random vectors to show that the outputs of multiple runs of hard-SVM on independent samples are typically concentrated around their mean.
@@
Let $\bm{x}_{1},\dots,\bm{x}_{k}\in\mathbb{R}^{d}$ be i.i.d random variables with mean $\mu$ and $\|\bm{x}_{i}-\mu\|_{\infty}\leq C$. Let $\bm{Z}=\frac{1}{k}\sum_{i=1}^{k}\bm{x}_{i}$. For all $t>0$,
$\Pr[\|\bm{Z}-\mu\|_{1}\geq t]\leq 2de^{\frac{-kt^{2}}{2d^{2}C^{2}}}.$
@@
We apply Hoeffding’s inequality to each coordinate and take the union bound. By Hoeffding’s inequality, for every $j\in[d]$, we have
@@
    The set $R_{x}\coloneqq\left\{\mathrm{round}_{\alpha}(y)\ |\ y\in\mathbb{S}^{d-1}\text{and }\|x-y\|_{2}\leq\beta\right\}$ has size at most $d$.
@@
We will construct a list-replicable learner that always outputs a hypothesis of the form $\overline{h}_{w}$ for some $w\in\mathbb{S}^{d-1}$.
@@
7:  output the hypothesis $\overline{h}_{\tilde{\bm{z}}}$.
@@
Let $\bm{\mathcal{A}}$ and $\bm{w}$ be as in Algorithm 1. For every $\epsilon,\delta\in(0,1)$ and $k\in\mathbb{N}$, there exists $n_{0}\coloneqq n_{0}(\epsilon,\delta,k)\in\mathbb{N}$ such that for every distribution $\mathcal{D}$ realizable by $\mathcal{H}_{\gamma}^{d}$, we have
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}\left[\|\bm{w}\|_{2}<\frac{\gamma}{2}\right]\leq\frac{\delta}{4}$ (4)
@@
$\Pr_{\mathbf{S}\sim\mathcal{D}^{kn_{0}}}[\mathcal{L}_{\mathcal{D}}(\bm{\mathcal{A}}(\mathbf{S}))\geq\epsilon]\leq\frac{\delta}{4}.$ (5)

### v4

@@ 2.2 Concluding remarks and open problems
For total concept classes, list replicability, shared randomness replicability, and (approximate) DP-learnability are now known to coincide through the combinatorial framework of the Littlestone dimension. In contrast, the situation for partial classes, as illustrated by the results of this paper, is more intricate and less understood.
The “DP-learnability to Shared-randomness replicability” reduction from [BGH+23] extends to the partial setting. Moreover, [FHM+24] recently showed that for partial classes, DP-learnability implies a finite Littlestone dimension, and [KKMV23, Lemma 8] shows that even in the partial setting, list replicability implies shared-randomness replicability.
Theorem 2.10 ([BGH+23, FHM+24, KKMV23]).
Let $\mathcal{C}\subseteq\left\{\pm 1,\star\right\}^{\mathcal{X}}$ be a partial concept class.
  • If $\mathcal{C}$ is DP-learnable, then $\mathcal{C}$ is shared-randomness replicable.
  • If $\mathcal{C}$ is DP-learnable, then $\operatorname{Ldim}(\mathcal{C})<\infty$.
  • If $\mathcal{C}$ is list replicable, then $\mathcal{C}$ is shared-randomness replicable.
On the other hand, our main theorem shows that for partial concepts, list replicability does not follow from bounded Littlestone dimension, shared-randomness replicability, DP-learnability, or even pure DP-learnability. To our knowledge, no further relationships among DP-learnability, shared-randomness replicability, Littlestone dimension, and list replicability are currently known for partial concept classes.
@@
    Is there a simple combinatorial notion of dimension that characterizes list replicability?
@@
    $\operatorname{VCdim}(\mathcal{C})\leq\widetilde{\operatorname{LR}}(\mathcal{C})\leq\operatorname{sign-rank}(\mathcal{C})?$
    Is it possible to upper-bound $\widetilde{\operatorname{LR}}(\mathcal{C})$ by a function of $\operatorname{VCdim}(\mathcal{C})$?
@@
    Question 4 in [AHHM21] also asks if every disambiguation of $\mathcal{H}_{\gamma}^{d}$ satisfies $\operatorname{VCdim}=\omega(1)$. Chornomaz, Moran, and Waknine explored this problem using a topological approach, but the question remains open  [CMW25].
3 Proof of Theorem 2.1
@@ 3.1 The lower bound
By the definition of list replicability, for any $\delta>0$, there is an integer $n$ so that for any realizable distribution $\mu$, there exists a list of hypotheses $\{h_{1},\dots,h_{L}\}$ with
$\operatorname{Pr}_{\mathbf{S}\sim\mu^{n}}[\bm{\mathcal{A}}(\mathbf{S})\in\{h_{1},\dots,h_{L}\}]\geq 1-\delta\text{ and }\operatorname{loss}_{\mu}(h_{i})\leq\epsilon\text{ for all }i\in[L].$
Now pick any $\alpha>0$ and $\epsilon^{\prime}\in(\epsilon,1/2)$. By taking $\delta$ sufficiently small, for any distribution $\mu$, we can choose a hypothesis $h_{\mu}\in\{h_{1},\dots,h_{L}\}$ such that
$\Pr_{\mathbf{S}\sim\mu^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{\mu}]>\frac{1}{L+\alpha}\text{ and }\operatorname{loss}_{\mu}(h_{\mu})<\epsilon^{\prime}.$ (4)
We will focus on a certain collection of realizable distributions $\mu$ on $\mathbb{S}^{d-1}\times\{\pm 1\}$. For any $w\in\mathbb{S}^{d-1}$, take $\mu_{w}$ to be the uniform distribution on the set $\{(x,c_{w}(x))\ |\ x\in{\mathrm{supp}}(c_{w})\}$. These distributions are, by definition, realizable. Hence, for each $\mu_{w}$, we can choose some particular hypothesis $h_{\mu_{w}}$ that satisfies the conditions in (4). Collect these hypotheses in a set $T$, that is to say
$T\coloneqq\{h_{\mu_{w}}\ |\ w\in\mathbb{S}^{d-1}\}.$
For each $h\in T$, define the set $U_{h}\subset\mathbb{S}^{d-1}$ as
$U_{h}\coloneqq\left\{w\in\mathbb{S}^{d-1}\ |\ \Pr_{\mathbf{S}\sim\mu_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h]>\frac{1}{L+\alpha}\text{ and }\operatorname{loss}_{\mu_{w}}(h)<\epsilon^{\prime}\right\}.$
@@
The family $\{U_{h}\}_{h\in T}$ forms an antipodal-free open cover of $\mathbb{S}^{d-1}$.
@@
The fact that any set $U_{h}$ is antipodal-free follows from the accuracy constraint $\operatorname{loss}_{\mu_{w}}(h)<\epsilon^{\prime}$. Indeed, for any $w\in\mathbb{S}^{d-1}$, the concepts $c_{w}$ and $c_{-w}$ have identical support, on which they disagree at every point. Thus the population loss of any hypothesis $h$ satisfies the equation
$\operatorname{loss}_{\mu_{w}}(h)+\operatorname{loss}_{\mu_{-w}}(h)=1.$
For any $w\in U_{h}$, we have that $\operatorname{loss}_{\mu_{w}}(h)<\epsilon^{\prime}<1/2$, whereby $w$ and w-w cannot both be in $U_{h}$.
Next, each set $U_{h}$ is open because both $\Pr_{\bm{S}\sim\mu_{w}^{n}}[\bm{\mathcal{A}}(\bm{S})=h]$ and $\operatorname{loss}_{\mu_{w}}(h)$ are continuous in $w$. Lastly, the family $\{U_{h}\}_{h\in T}$ covers $\mathbb{S}^{d-1}$ because, for any $w\in\mathbb{S}^{d-1}$, the set $U_{h_{\mu_{w}}}$ contains $w$ by construction. ∎
Now note that the antipodal-free open cover $\{U_{h}\}_{h\in T}$ admits a finite subcover by the compactness of the unit sphere. Applying Theorem 3.1 to such a finite subcover guarantees that some $w\in\mathbb{S}^{d-1}$ is contained in at least $t\coloneqq\lceil\frac{d}{2}+1\rceil$ sets $U_{h_{1}},U_{h_{2}},\dots,U_{h_{t}}$. Unpacking definitions reveals that the distribution $\mu_{w}$ has the property
$\Pr_{\mathbf{S}\sim\mu_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]>\frac{1}{L+\alpha}$
for $t$ distinct hypotheses $h_{i}\in T$. Because these $h_{i}$ are distinct, the events $\bm{\mathcal{A}}(\mathbf{S})=h_{i}]$ are disjoint, and therefore
$1\geq\operatorname{Pr}_{\mathbf{S}\sim\mu_{w}^{n}}\bigcup_{i=1}^{t}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]=\sum_{i=1}^{t}\operatorname{Pr}_{\mathbf{S}\sim\mu_{w}^{n}}[\bm{\mathcal{A}}(\mathbf{S})=h_{i}]>\frac{t}{L+\alpha}.$
@@ 3.2 The upper bound
To prove the upper bound, we design a list replicable learning algorithm $\bm{\mathcal{A}}$ that learns $\mathcal{H}_{\gamma}^{d}$ with list size $d$ independent of $\epsilon>0$. Given $w\in\mathbb{S}^{d-1}$, let $\overline{c}_{w}:\mathbb{S}^{d-1}\to\{\pm 1\}$ denote the total concept corresponding to the closed half-space defined by $w$.
$\overline{c}_{w}(x)\coloneqq\begin{cases}1&\text{if }{\left\langle w,x\right\rangle}\geq 0\\ -1&\text{if }{\left\langle w,x\right\rangle}<0\end{cases}.$
@@
[…]
@@
Furthermore, for any distribution $\mu$ over $\mathbb{S}^{d-1}\times\{\pm 1\}$, we call $w$ a $(\gamma,\epsilon)$-separator for $\mu$ if
$\Pr_{(x,y)\sim\mu}[y{\left\langle x,w\right\rangle}<\gamma]\leq\epsilon.$
When learning $\mathcal{H}_{\gamma}^{d}$ in the realizable setting, for any sample set $S$ drawn from a realizable distribution $\mu$, there is some $w$ that $\gamma$-separates $S$. Therefore, the hard-SVM will be feasible and return a vector $w$ that $\gamma$-separates $S$.
The following theorem, due to [STBWA98], says that if we take a sufficiently large sample $\bm{S}$ and compute a good separator $w$ for it using hard-SVM, then with high probability, $w$ will also be a good separator for $\mu$.
@@
For all $\epsilon,\delta>0$, there exists $n\coloneqq n(\epsilon,\delta)$ such that the following holds. Let $\mu$ be any distribution over $\mathbb{S}^{d-1}\times\left\{\pm 1\right\}$.
$\Pr_{\bm{S}\sim\mu^{n}}\left[\text{Every $w\in\mathbb{S}^{d-1}$ that $\gamma$-separates $\bm{S}$ also $\left(\frac{\gamma}{2},\epsilon\right)$-separates $\mu$}\right]\geq 1-\delta.$
@@
To prove Theorem 3.4, one can apply [STBWA98, Theorem 3.5] to show that, with probability at least $1-\delta$, both $h_{1}(x)\coloneqq\operatorname{sgn}({\left\langle x,w\right\rangle}+\frac{\gamma}{2})$ and $h_{2}(x)\coloneqq\operatorname{sgn}({\left\langle x,w\right\rangle}-\frac{\gamma}{2})$ have loss at most $\frac{\epsilon}{2}$, in which case $w$ is a $\left(\frac{\gamma}{2},\epsilon\right)$-separator for $\mu$.
@@
First, we prove a simple concentration result for sums of i.i.d. random vectors to show that the outputs of multiple runs of hard-SVM on independent samples are typically concentrated around their mean.
@@
[…]
@@
We apply Hoeffding’s inequality  [^4] Let $c\in\mathbb{R}$ and let $\bm{x}_{1},\ldots,\bm{x}_{n}$ be independent random variables with $\bm{x}_{i}\in[-c,c]$ and $\mathbb{E}[\bm{x}_{i}]=0$. For any $t>0$, $\Pr\left[\left|\sum_{i=1}^{n}\bm{x}_{i}\right|\geq t\right]\leq 2e^{-\frac{t^{2}}{2c}}.$ to each coordinate and take the union bound. By Hoeffding’s inequality, for every $j\in[d]$, we have
@@
[…]
@@
We will construct a list-replicable learner that always outputs a hypothesis of the form $\overline{c}_{w}$ for some $w\in\mathbb{S}^{d-1}$.
@@
7:output the hypothesis $\overline{c}_{\tilde{\bm{z}}}$.
@@
Let $\bm{\mathcal{A}}$ and $\bm{w}$ be as in Algorithm 1. For every $\epsilon,\delta\in(0,1)$ and $k\in\mathbb{N}$, there exists $n_{0}\coloneqq n_{0}(\epsilon,\delta,k)\in\mathbb{N}$ such that for every distribution $\mu$ realizable by $\mathcal{H}_{\gamma}^{d}$, we have
[…]
@@
$\Pr_{\mathbf{S}\sim\mu^{kn_{0}}}[\operatorname{loss}_{\mu}(\bm{\mathcal{A}}(\mathbf{S}))\geq\epsilon]\leq\frac{\delta}{4}.$ (6)